Calculate the Zeff for the valence electrons in the atoms Li to Ne using (a) the assumption that o equals the number of inner-shell electrons and (b) Slater's rules. Plot both sets of results on the same graph and discuss.
Answer #2

Atom

Atomic Number

Electronic configuration

Li

Be

B

C

N

O

F

Ne

3

4

5

6

7

8

9

10

1s2 2s1

1s2 2s2

1s2 2s2p1

1s2 2s2p2

1s2 2s2p3

1s2 2s2p4

1s2 2s2p5

1s2 2s2p6

a. Calculating  Zeff when σ is the number of inner shell electrons ie electrons in the n-1 shell ie only taking the contribution from the 1s2 electrons

Zeff of Li valence electron = 3 - 2 = 1

Zeff of Be valence electron = 4 - 2 = 2

Zeff of B valence electron = 5 - 2 = 3

Zeff of C valence electron = 6 - 2 = 4

Zeff of N valence electron = 7 - 2 = 5

Zeff of O valence electron= 8 - 2 = 6

Zeff of F valence electron = 9 - 2 = 7

Zeff of Ne valence electron = 10 - 2 = 8

b. Effective nuclear charge Zeff is given as: Zeff = Z - σ

where Z is the atomic number or the number of protons and σ is the screening or the shielding constant

When we calculate the Zeff of the valence electrons of the atoms from Li to Ne as per the Slater's rules:

Zeff of Li valence electron= 3 - σ = 3 - (2 x 0.35) = 3 - 0.70 = 2.30

Zeff of Be valence electron = 4 - (1 x 0.35 + 2 x 0.85) = 1.95

Zeff of B valence electron = 5 - (2 x 0.35 + 2 x 0.85) = 2.60

Zeff of C valence electron = 6 - (3 x 0.35 + 2 x 0.85) = 3.25

Zeff of N valence electron = 7 - (4 x 0.35 + 2 x 0.85) = 3.9

Zeff of O valence electron = 8 - (5 x 0.35 + 2 x 0.85) = 4.55

Zeff of F valence electron = 9 - (6 x 0.35 + 2 x 0.85) = 5.2

Zeff of Ne valence electron = 10 - (7 x 0.35 + 2 x 0.85) = 5.85

Zeff graph is plotted with both the Zeff calculated values as given below

y-axis gives Zeff and x-axis gives the atomic number of the elements from Be to Li as given in the table above. Series1 gives the plot for Zeff considering only the inner shell electrons and Series2 gives the plot from calculations using the Slaters rule.
Answer #1

Formula: Zeff = Z -

Part (a).

Li: Zeff = 3 - 2 = 1

Be: Zeff = 4 - 2 = 2

B: Zeff = 5 - 2 = 3

C: Zeff = 6 - 2 = 4

N: Zeff = 7 - 2 = 5

O: Zeff = 8 - 2 = 6

F: Zeff = 9 - 2 = 7

Ne: Zeff = 10 - 2 = 8

Part (b).

Li: Zeff = 3 - (2*0.85) = 1.3

Be: Zeff = 4 - (1*0.35 + 2*0.85) = 1.95

B: Zeff = 5 - (2*0.35 + 2*0.85) = 2.6

C: Zeff = 6 - (3*0.35 + 2*0.85) = 3.25

N: Zeff = 7 - (4*0.35 + 2*0.85) = 3.9

O: Zeff = 8 - (5*0.35 + 2*0.85) = 4.55

F: Zeff = 9 - (6*0.35 + 2*0.85) = 5.2

Ne: Zeff = 10 - (7*0.35 + 2*0.85) = 5.85